# Strand, London WC2R 2LS, UK

### v1

@@ 2.2 On-shell action and conformal thermodynamics
This can be immediately recognized to be in the desired form (22), after identifying $\tau\rightarrow e^{\boldsymbol{\omega}}u/\sqrt{f(e^{{\boldsymbol{\omega}}}\mathfrak{r})}$.
As usual, the Euclidean time is periodically identified as $\tau\sim\tau+\beta$. Following the classic work of York York:1986it , the canonical ensemble is defined by fixing the proper boundary temperature $\beta_{\partial\mathcal{M}}$ along with any other charges. For (extended) conformal boundary conditions, $\beta_{\partial\mathcal{M}}$ can in principle fluctuate, so instead we define the (dimensionless) inverse temperature
@@
In other words, the gravitational partition function in in the saddle point approximation, with the saddle obeying ECBC, is given by
@@
Given $\mathcal{Z}_{\text{grav}}[\tilde{\beta},\mathcal{P},Q]$, we follow the Gibbons-Hawking prescription Gibbons:1976ue to compute the entropy, energy and specific heat of this conformal canonical ensemble as
@@
Before actually computing thermodynamic quantities in our setup, let us briefly comment on some expectations coming from thermal effective field theory. As emphasized in Banihashemi:2024yye ; Banihashemi:2025qqi , one can organize the high-temperature limit of the conformal thermodynamics (for GR) in terms of an effective action of a thermal CFT in one dimension lower.
@@
Assuming that the theory is trivially gapped at finite temperature, in the $\tilde{\beta}\rightarrow 0$ limit, one can dimensionally reduce over the thermal circle, yielding an effective action of the form Banerjee:2012iz ; Benjamin:2023qsc  [^7] Note that $c_{0}$ is proportional to the thermal one-point function of the stress tensor $\langle T^{\mu\nu}(t,\vec{x})\rangle_{\tilde{\beta}}$ on $\mathbb{R}^{d-1}$, or alternatively to the Casimir energy on $\mathbb{S}^{1}\times\mathbb{R}^{d-2}$ Benjamin:2023qsc . Similarly, $c_{1}$ is related to the subleading contribution of the Casimir energy when the latter is compactified to $\mathbb{S}^{1}\times\mathbb{S}^{d-2}$ Allameh:2024qqp .
@@
where $R_{\gamma}$ is the Ricci scalar of the ‘spatial’ metric $\gamma$, e.g., see section $1.3$ of ref. Banihashemi:2024yye . From the thermal effective action, we can compute the thermal entropy as usual,
@@
where we have introduced $\Sigma_{k,d-1}$ as the volume and substituted $R_{\gamma}=\frac{k\,(d-1)(d-2)}{\mathfrak{r}^{2}}$ for the constant curvature metrics of curvature radius $\mathfrak{r}$ in eq. (16). Moreover, it was conjectured in Allameh:2024qqp that the subleading coefficient $c_{1}>0$. This conjecture is supported by examining free field theories and thermal CFTs with an Einstein gravity dual. Nevertheless, it was observed in Banihashemi:2024yye ; Banihashemi:2025qqi that thermodynamics of CBC consistently violate this bound. In turn, this violation was interpreted as a sign that gravity does not completely decouple in this holographic setup. Instead, it was argued that the boundary conformal mode fluctuates and must be integrated over in a corresponding CFT path integral.
@@ 3 Euclidean Conformal Thermodynamics
Note that at the boundary $y=r_{h}/e^{\boldsymbol{\omega}}\mathfrak{r}$. In principle, the relations defining the (E)CBC (27) could be inverted to express $r_{h}$ and $e^{\boldsymbol{\omega}}\mathfrak{r}$ as functions of the boundary data $\{\tilde{\beta},\mathcal{P},Q\}$. However this becomes analytically intractable even in pure Einstein gravity (with the exception of the $d=2$ case, that can be solved exactly Anninos:2024wpy ).
One alternative is to proceed by solving the full problem numerically. This was done, for instance, in Anninos:2024wpy for the positive cosmological constant case in $d=3$, and in Banihashemi:2025qqi for charged black holes in $d=3$. Here we will be mostly interested in the high-temperature behavior of the partition function, so, as we will show, we only need to invert the expressions perturbatively in this limit.
@@ 3.1 Conformal Thermodynamics in Einstein Gravity
In this subsection, we briefly review Euclidean conformal thermodynamics in Einstein gravity with an emphasis on the conformal entropy, to set the stage for our subsequent EMGB analysis. This problem has been analyzed previously in Anninos:2023epi ; Anninos:2024wpy ; Banihashemi:2024yye ; Banihashemi:2025qqi ; Liu:2024ymn .
Consider pure Einstein gravity with negative cosmological constant in $d+1$ dimensions. If we only consider stationary solutions, by Birkhoff’s theorem we know that the line element (20) with $q=0$ covers the full solution space of the problem, up to solutions that differ in their boundary Weyl factor. [^8] This is because the bulk Killing vector fields act as conformal Killing vectors at the boundary. Hence, Birkhoff’s theorem only guarantees the uniqueness for the solution in the bulk, up to possibly a finite number of isolated solutions with non-trivial Weyl factor. We will return to this point in Section 4, when we discuss dynamics in Lorentzian signature. For the spherical black hole case (i.e., k=+1k=+1), [^9] The results readily generalize to all maximally symmetric horizon topologies, as we discuss in the EMGB analysis below. the CBC (3) for the line element (20) read
@@
One way of explaining this fact is that black hole entropy can be computed using Wald’s formula Wald:1993nt (or the generalizations in PhysRevLett.70.3684 ; Jacobson:1993vj ; Jacobson:1994qe ; Dong:2013qoa ; Camps:2013zua ). Alternatively, it is possible to derive this universal formula only assuming the existence of a local, covariant boundary term Carlip:1993sa ; Banados:1993qp . We review this derivation in appendix F. We will later generalize these arguments for general Lovelock theories.
@@
where $\Omega_{d-1}=\tfrac{2\pi^{d/2}}{\Gamma({d}/{2})}$ is the area of a unit (d1)(d{-}1)-sphere. Here, we see that this conformal entropy agrees with the structure of (30), which corresponds to the thermal entropy of a $d$-dimensional CFT . Subleading corrections of this high-temperature asymptotic expansion can also be matched Banihashemi:2024yye ; Banihashemi:2025qqi . The results for positive and vanishing cosmological constant can be easily obtained from this one. For the flat space limit, we just take the dimensionless combination $\mathcal{K}\ell\to\infty$, while to obtain the positive cosmological constant answer, we analytically continue $\ell\rightarrow i\ell$. [^11] Throughout, we choose the principal branch in defining $\mathcal{N}_{\text{dof}}$ as an analytic function of $\ell$. In summary, in all cases we obtain
@@
$\mathcal{N}_{\text{dof}}^{(\text{GR})}(\mathcal{K})=\begin{cases}\dfrac{\Omega_{d-1}\ell^{d-1}}{4G_{N}}\left(\dfrac{4\pi}{d^{2}}\right)^{d-1}\left(\sqrt{\mathcal{K}^{2}\ell^{2}+d^{2}}-\mathcal{K}\ell\right)^{d-1}&\quad\Lambda>0\ ,\\[8.61108pt] \dfrac{\Omega_{d-1}\left(2\pi\right)^{d-1}}{4G_{N}\mathcal{K}^{d-1}}\ &\quad\Lambda=0\ ,\\[8.61108pt] \dfrac{\Omega_{d-1}\ell^{d-1}}{4G_{N}}\left(\dfrac{4\pi}{d^{2}}\right)^{d-1}\left(\mathcal{K}\ell-\sqrt{\mathcal{K}^{2}\ell^{2}-d^{2}}\right)^{d-1}\ &\quad\Lambda<0\ .\end{cases}$ (41)
@@
An interesting observation is that $\mathcal{N}^{(\text{GR})}_{\text{dof}}(\mathcal{K})$ is a monotonic function. For negative cosmological constant, it decreases from the conformal boundary of AdS, where $\mathcal{K}\ell=d$, all the way to the black hole horizon, where $\mathcal{K}\ell\to\infty$. Of course, when $\mathcal{K}\ell=d$, we recover the standard AdS/CFT result. For instance, when $d=2$, we obtain that $\mathcal{N}_{\text{dof}}(\mathcal{K})$ at the boundary is the Brown-Henneaux central charge. This monotonic behavior is somewhat reminiscent of the holographic $c$-functions Freedman:1999gp ; Myers:2010tj ; Myers:2012ed . It would be interesting to have a microscopic interpretation of this number of degrees of freedom [^12] A specific proposal for $d=2$ is given in Allameh:2025gsa . – see section 5 for further discussion.
@@
Notice that in contrast to the conformal case discussed previously, the entropy here approaches to a constant value at high temperatures. Further, the leading correction away from this constant goes to zero in a way that does not depend on the spacetime dimension. In AdS/CFT duality, these finite boundaries appear to have an interpretation as irrelevant deformations of the boundary theory Freidel:2008sh ; McGough:2016lol ; Zamolodchikov:2004ce ; Smirnov:2016lqw ; Taylor:2018xcy ; Hartman:2018tkw . For instance, in $d=2$ this corresponds to the $T\bar{T}$ deformation of Zamolodchikov Zamolodchikov:2004ce .
@@ 3.2 Conformal Thermodynamics in Einstein-Gauss-Bonnet Gravity
$\mathcal{S}^{\text{(EMGB)}}=\frac{\Sigma_{k,d-1}r_{h}^{d-1}}{4G_{N}}\left(1+2k\frac{d-1}{d-3}\frac{{\alpha}}{r_{h}^{2}}\right)\ ,$ (45)
which is the well-known result derived in Jacobson:1994qe that corresponds to the Wald entropy for GB black holes Wald:1993nt . This bulk formula holds independently of the boundary conditions, the sign of the cosmological constant and the matter content of the theory. The derivation only requires the existence of a local, covariant boundary term. We review this derivation in appendix F.
@@ 3.2.1 EGB AdS Black Hole
as can be seen by taking the $r\rightarrow\infty$ limit of eq. (46). The distinction between $\ell$ and $\ell_{\textrm{\tiny AdS}}$ will become important when trying to compare to the pure Einstein gravity results.
@@
Indeed, the non-perturbative GB contribution takes a finite value at the AdS boundary. This value coincides with the one expected from AdS/CFT, which for finite GB coupling takes the form Buchel:2009sk
@@
As we increase ${\mathcal{P}}\ell_{\textrm{\tiny AdS}}$, the boundary is driven into the interior of the bulk, and $\mathcal{N}_{\text{dof}}(\mathcal{P}\ell_{\textrm{\tiny AdS}})$ decreases monotonically towards the black hole horizon, which has ${\mathcal{P}}\ell_{\textrm{\tiny AdS}}\to\infty$. For ${\mathcal{P}}\ell_{\textrm{\tiny AdS}}$ away from the boundary, $\mathcal{N}_{\text{dof}}(\mathcal{P}\ell_{\textrm{\tiny AdS}})$ becomes indistinguishable from the Einstein gravity result. As in the Einstein case, this behavior is reminiscent of that for holographic $c$-functions Freedman:1999gp ; Myers:2010tj ; Myers:2012ed .
@@ 3.2.2 EGB dS Black Hole
Similarly, we can study the problem in EGB with a positive cosmological constant. Given the existence of both black hole and cosmological horizons, in this case, there are different types of patches that can be considered. Ref. Anninos:2024wpy presents an exhaustive analysis of spherically symmetric static configurations in Einstein gravity. Here, we will just restrict ourselves to what are called cosmic patches, that is the bulk region in between the boundary and the cosmological horizon.
@@
Performing a similar analysis as in the AdS case above, one arrives at expressions for the conformal entropy that are related to the AdS ones via a simple analytic continuation $\ell\rightarrow i\ell$ in eq. (57), recall footnote 11. More specifically, the dS conformal entropy reads
@@
We plot eq. (65) in figure 2. We observe that as in Einstein gravity, $\mathcal{N}_{\text{dof}}(\mathcal{P}\ell_{\textrm{\tiny dS}})$ is a monotonically decreasing function of ${\mathcal{P}}\ell_{\textrm{\tiny dS}}$. As opposed to the AdS case, the perturbative expansion doesn’t break down, and it is a good approximation to the non-perturbative result for small $\tilde{\alpha}$, for all allowed values of ${\mathcal{P}}\ell_{\textrm{\tiny dS}}$.
@@ 3.2.3 Flat EGB Black Hole
Taking the large-$\mathcal{P}$$\ell$ limit of, e.g., the conformal entropy for the AdS black hole (57), we obtain [^17] Note the solutions with maximally symmetric horizons for $k=0$ and –1 are not conventional black holes, but correspond to what are known as flat space cosmologies, see e.g., ref Banihashemi:2025qqi .
@@
Note that the leading term in the high-temperature expansion does not receive a correction in the GB coupling (recall footnote 14). The first non-trivial correction comes at order $\mathcal{O}(\tilde{\beta}^{3-d})$. This is unlike the AdS and dS cases, and might be a distinct feature of quantum gravity in flat space. The same result can be obtained directly from the flat space black hole, but then the GB coupling needs to scale with $G_{N}$ – see appendix D.

### v2

@@ 2.2 On-shell action and conformal thermodynamics
This can be immediately recognized to be in the desired form (22), after identifying $d\tau\rightarrow e^{\boldsymbol{\omega}}du/\sqrt{f(e^{{\boldsymbol{\omega}}}\mathfrak{r})}$.
As usual, the Euclidean time is periodically identified as $\tau\sim\tau+\beta$. Following the classic work of York York (1986), the canonical ensemble is defined by fixing the proper boundary temperature $\beta_{\partial\mathcal{M}}$ along with any other charges. For (extended) conformal boundary conditions, $\beta_{\partial\mathcal{M}}$ can in principle fluctuate, so instead we define the (dimensionless) inverse temperature
@@
In other words, the gravitational partition function in the saddle point approximation, with the saddle obeying ECBC, is given by
@@
Given $\mathcal{Z}_{\text{grav}}[\tilde{\beta},\mathcal{P},Q]$, we follow the Gibbons-Hawking prescription Gibbons and Hawking (1977) to compute the entropy, energy and specific heat of this conformal canonical ensemble as
@@
Before actually computing thermodynamic quantities in our setup, let us briefly comment on some expectations coming from thermal effective field theory. As emphasized in Banihashemi et al. (2025a, b), one can organize the high-temperature limit of the conformal thermodynamics (for GR) in terms of an effective action of a thermal CFT in one dimension lower.
@@
Assuming that the theory is trivially gapped at finite temperature, in the $\tilde{\beta}\rightarrow 0$ limit, one can dimensionally reduce over the thermal circle, yielding an effective action of the form Banerjee et al. (2012); Benjamin et al. (2024) [^7] Note that $c_{0}$ is proportional to the thermal one-point function of the stress tensor $\langle T^{\mu\nu}(t,\vec{x})\rangle_{\tilde{\beta}}$ on $\mathbb{R}^{d-1}$, or alternatively to the Casimir energy on $\mathbb{S}^{1}\times\mathbb{R}^{d-2}$ Benjamin et al. (2024). Similarly, $c_{1}$ is related to the subleading contribution of the Casimir energy when the latter is compactified to $\mathbb{S}^{1}\times\mathbb{S}^{d-2}$ Allameh and Shaghoulian (2025a).
@@
where $R_{\gamma}$ is the Ricci scalar of the ‘spatial’ metric $\gamma$, e.g., see section $1.3$ of ref. Banihashemi et al. (2025a). From the thermal effective action, we can compute the thermal entropy as usual,
@@
where we have introduced $\Sigma_{k,d-1}$ as the volume and substituted $R_{\gamma}=\frac{k\,(d-1)(d-2)}{\mathfrak{r}^{2}}$ for the constant curvature metrics of curvature radius $\mathfrak{r}$ in eq. (16). Moreover, it was conjectured in Allameh and Shaghoulian (2025a) that the subleading coefficient $c_{1}>0$. This conjecture is supported by examining free field theories and thermal CFTs with an Einstein gravity dual. Nevertheless, it was observed in Banihashemi et al. (2025a, b) that thermodynamics of CBC consistently violate this bound. In turn, this violation was interpreted as a sign that gravity does not completely decouple in this holographic setup. Instead, it was argued that the boundary conformal mode fluctuates and must be integrated over in a corresponding CFT path integral.
@@ 3 Euclidean Conformal Thermodynamics
Note that at the boundary $y=r_{h}/e^{\boldsymbol{\omega}}\mathfrak{r}$. In principle, the relations defining the (E)CBC (27) could be inverted to express $r_{h}$ and $e^{\boldsymbol{\omega}}\mathfrak{r}$ as functions of the boundary data $\{\tilde{\beta},\mathcal{P},Q\}$. However this becomes analytically intractable even in pure Einstein gravity (with the exception of the $d=2$ case, that can be solved exactly Anninos et al. (2024)).
One alternative is to proceed by solving the full problem numerically. This was done, for instance, in Anninos et al. (2024) for the positive cosmological constant case in $d=3$, and in Banihashemi et al. (2025b) for charged black holes in $d=3$. Here we will be mostly interested in the high-temperature behavior of the partition function, so, as we will show, we only need to invert the expressions perturbatively in this limit.
@@ 3.1 Conformal Thermodynamics in Einstein Gravity
In this subsection, we briefly review Euclidean conformal thermodynamics in Einstein gravity with an emphasis on the conformal entropy, to set the stage for our subsequent EMGB analysis. This problem has been analyzed previously in Anninos et al. (2023a, 2024); Banihashemi et al. (2025a, b); Liu et al. (2024).
Consider pure Einstein gravity with negative cosmological constant in $d+1$ dimensions. [^8] Note that in Einstein gravity, in contrast to more general Lovelock theories, the length scale $\ell$ appearing in the action (9) consides with the (A)dS curvature radius $\ell_{\textrm{\tiny\text{(A)dS}}}$ and thus we do not distinguish between the two in this subsection. If we only consider stationary solutions, by Birkhoff’s theorem we know that the line element (20) with $q=0$ covers the full solution space of the problem, up to solutions that differ in their boundary Weyl factor. [^9] This is because the bulk Killing vector fields act as conformal Killing vectors at the boundary. Hence, Birkhoff’s theorem only guarantees the uniqueness for the solution in the bulk, up to possibly a finite number of isolated solutions with non-trivial Weyl factor. We will return to this point in Section 4, when we discuss dynamics in Lorentzian signature. For the spherical black hole case (i.e., k=+1k=+1), [^10] The results readily generalize to all maximally symmetric horizon topologies, as we discuss in the EMGB analysis below. the CBC (3) for the line element (20) read
@@
One way of explaining this fact is that black hole entropy can be computed using Wald’s formula Wald (1993) (or the generalizations in Jacobson and Myers (1993); Jacobson et al. (1994b, a); Dong (2014); Camps (2014)). Alternatively, it is possible to derive this universal formula only assuming the existence of a local, covariant boundary term Carlip and Teitelboim (1995); Banados et al. (1994). We review this derivation in appendix F. We will later generalize these arguments for general Lovelock theories.
@@
where $\Omega_{d-1}=\tfrac{2\pi^{d/2}}{\Gamma({d}/{2})}$ is the area of a unit (d1)(d{-}1)-sphere. Here, we see that this conformal entropy agrees with the structure of (30), which corresponds to the thermal entropy of a $d$-dimensional CFT . Subleading corrections of this high-temperature asymptotic expansion can also be matched Banihashemi et al. (2025a, b). The results for positive and vanishing cosmological constant can be easily obtained from this one. For the flat space limit, we just take the dimensionless combination $\mathcal{K}\ell\to\infty$, while to obtain the positive cosmological constant answer, we analytically continue $\ell\rightarrow i\ell$. [^12] Throughout, we choose the principal branch in defining $\mathcal{N}_{\text{dof}}$ as an analytic function of $\ell$. In summary, in all cases we obtain
@@
An interesting observation is that $\mathcal{N}^{(\text{GR})}_{\text{dof}}(\mathcal{K})$ is a monotonic function. For negative cosmological constant, it decreases from the conformal boundary of AdS, where $\mathcal{K}\ell=d$, all the way to the black hole horizon, where $\mathcal{K}\ell\to\infty$. Of course, when $\mathcal{K}\ell=d$, we recover the standard AdS/CFT result. For instance, when $d=2$, we obtain that $\mathcal{N}_{\text{dof}}(\mathcal{K})$ at the boundary is the Brown-Henneaux central charge. This monotonic behavior is somewhat reminiscent of the holographic $c$-functions Freedman et al. (1999); Myers and Sinha (2011); Myers and Singh (2012). It would be interesting to have a microscopic interpretation of this number of degrees of freedom [^13] A specific proposal for $d=2$ is given in Allameh and Shaghoulian (2025b). – see section 5 for further discussion.
@@
Notice that in contrast to the conformal case discussed previously, the entropy here approaches to a constant value at high temperatures. Further, the leading correction away from this constant goes to zero in a way that does not depend on the spacetime dimension. In AdS/CFT duality, these finite boundaries appear to have an interpretation as irrelevant deformations of the boundary theory Freidel (2008); McGough et al. (2018); Zamolodchikov (2004); Smirnov and Zamolodchikov (2017); Taylor (2023); Hartman et al. (2019). For instance, in $d=2$ this corresponds to the $T\bar{T}$ deformation of Zamolodchikov Zamolodchikov (2004).
@@ 3.2 Conformal Thermodynamics in Einstein-Gauss-Bonnet Gravity
$\mathcal{S}^{\text{(EMGB)}}_{\text{conf}}=\frac{\Sigma_{k,d-1}r_{h}^{d-1}}{4G_{N}}\left(1+2k\frac{d-1}{d-3}\frac{{\alpha}}{r_{h}^{2}}\right)\ ,$ (45)
which is the well-known result derived in Jacobson et al. (1994a) that corresponds to the Wald entropy for GB black holes Wald (1993). This bulk formula holds independently of the boundary conditions, the sign of the cosmological constant and the matter content of the theory. The derivation only requires the existence of a local, covariant boundary term. We review this derivation in appendix F.
@@ 3.2.1 EGB AdS Black Hole
as can be seen by taking the $r\rightarrow\infty$ limit of eq. (46). Note that we defined $\tilde{\alpha}:=\tfrac{\alpha}{\ell^{2}}$. The distinction between $\ell$ and $\ell_{\textrm{\tiny AdS}}$ will become important when trying to compare to the pure Einstein gravity results.
@@
Indeed, the non-perturbative GB contribution takes a finite value at the AdS boundary. This value coincides with the one expected from AdS/CFT, which for finite GB coupling takes the form Buchel et al. (2010)
@@
As we increase ${\mathcal{P}}\ell_{\textrm{\tiny AdS}}$, the boundary is driven into the interior of the bulk, and $\mathcal{N}_{\text{dof}}(\mathcal{P}\ell_{\textrm{\tiny AdS}})$ decreases monotonically towards the black hole horizon, which has ${\mathcal{P}}\ell_{\textrm{\tiny AdS}}\to\infty$. For ${\mathcal{P}}\ell_{\textrm{\tiny AdS}}$ away from the boundary, $\mathcal{N}_{\text{dof}}(\mathcal{P}\ell_{\textrm{\tiny AdS}})$ becomes indistinguishable from the Einstein gravity result. As in the Einstein case, this behavior is reminiscent of that for holographic $c$-functions Freedman et al. (1999); Myers and Sinha (2011); Myers and Singh (2012).
@@ 3.2.2 EGB dS Black Hole
Similarly, we can study the problem in EGB with a positive cosmological constant. Given the existence of both black hole and cosmological horizons, in this case, there are different types of patches that can be considered. Ref. Anninos et al. (2024) presents an exhaustive analysis of spherically symmetric static configurations in Einstein gravity. Here, we will just restrict ourselves to what are called cosmic patches, that is the bulk region in between the boundary and the cosmological horizon.
@@
Performing a similar analysis as in the AdS case above, one arrives at expressions for the conformal entropy that are related to the AdS ones via a simple analytic continuation $\ell\rightarrow i\ell$ in eq. (57), recall footnote 12. More specifically, the dS conformal entropy reads
@@
We plot eq. (65) in figure 2. We observe that as in Einstein gravity, $\mathcal{N}_{\text{dof}}(\mathcal{P}\ell_{\textrm{\tiny dS}})$ is a monotonically decreasing function of ${\mathcal{P}}\ell_{\textrm{\tiny dS}}$. As opposed to the AdS case, the perturbative expansion does not break down, and it is a good approximation to the non-perturbative result for small $\tilde{\alpha}$, for all allowed values of ${\mathcal{P}}\ell_{\textrm{\tiny dS}}$.
@@ 3.2.3 Flat EGB Black Hole
Taking the large-$\mathcal{P}$$\ell$ limit of, e.g., the conformal entropy for the AdS black hole (57), we obtain [^18] Note the solutions with maximally symmetric horizons for $k=0$ and –1 are not conventional black holes, but correspond to what are known as flat space cosmologies, see e.g., ref Banihashemi et al. (2025b).
@@
Note that the leading term in the high-temperature expansion does not receive a correction in the GB coupling (recall footnote 15). The first non-trivial correction comes at order $\mathcal{O}(\tilde{\beta}^{3-d})$. This is unlike the AdS and dS cases, and might be a distinct feature of quantum gravity in flat space. The same result can be obtained directly from the flat space black hole, but then the GB coupling needs to scale with $G_{N}$ – see appendix D.